gabrielle1:

Terri’s Ice Cream Parlor makes gourmet ice cream cones. The volume of each cone is 8 cubic inches. The height is 4 inches more than the radius of the cone’s opening. a.) Write a polynomial function that represents the volume of the ice cream cone in terms of the radius (the formula for the volume of a cone is 1 2 3 V r h   ). b.) What are the possible values of r ? c.) Find the dimensions of the cone.

Mercury:

re-formatting the question so it's a little easier to read: Terri’s Ice Cream Parlor makes gourmet ice cream cones. The volume of each cone is 8 cubic inches. The height is 4 inches more than the radius of the cone’s opening. a.) Write a polynomial function that represents the volume of the ice cream cone in terms of the radius (the formula for the volume of a cone is V = (1/3)pi * r^2 * h b.) What are the possible values of r? c) Find the dimensions of the cone.

Mercury:

a) "The height is 4 inches more than the radius of the cone’s opening." therefore, h = r + 4 "The volume of each cone is 8 cubic inches" therefore, V = 8 if you plug these into the volume formula, you'll have an equation in terms of r. solve for r.

Mercury:

b) once you've solved for r, consider which of your r-values is actually a viable solution to this problem. remember, r stands for the radius of the cone so it must be a positive number c) you already have r. since h = r + 4, simply add 4 to the radius to get the height. radius and height are the dimensions you'll report.